1.Timmy wants to buy a scooter and the price was P 3, 500. What he goes to the store a second time, he found that price was marked down by 20%. What is the new price?

2.Emily buys a toaster during the sale for 10% off. If Ellen pays P 3600, What was the Original price?


3.Zack has an old car. He wants to sell it for 60% off the current price. The market price is P 500, 000.00. How much money would he receive in exchange for the car if he were able to sell it at to that rate?
1
Expert's answer
2021-09-08T11:47:37-0400

1. Proportion:

3500 – 100 %

x – 20 %

"x = \frac{3500 \times 20}{100}=700"

New price = 3500 -700=2800

2. 100%-10%=90%

Proportion:

90% - 3600

100% - x

"x= \frac{100 \times 3600}{90}= 4000"

3. Proportion:

100 %-60%=40%

500000 – 100%

x – 40 %

"x = \frac{500000 \times 40 }{100}=200000"